# Focus On This, Not That! Steering LLMs With Adaptive Feature Specification

### v2

@@ 4 Experiments
In this section we empirically validate the effectiveness of FIT across a range of popular LLMs of varying sizes and on different NLP datasets, including classification and multi-choice question-answering tasks.
Before reporting the main results, we introduce the evaluation metric (focus accuracy) that we report, baselines, models, and training settings used throughout the experiments. In Section 4.1, we first verify that FIT performs well on the simpler SS dataset, a synthetic sentiment analysis dataset derived from SST-5 (Socher et al., 2013b). We then demonstrate in Section 4.2 that FIT generalises to more complex features and handles distribution shifts on the SMNLI dataset, a sub-sampled version of the MNLI dataset (Williams et al., 2018). Finally, in Section 4.3, we show that FIT has practical, real-world impact by effectively mitigating bias in the BBQ dataset (Parrish et al., 2022), where we further illustrate FIT’s ability to generalise to new features seen for the first time when performing inference.
Although the primary focus of FIT is on adaptively steering LLMs at inference time, we include an additional debiasing experiment for comparison on the BBQ dataset in Section A.8 for completeness. While bias mitigation is a valuable and natural application of FIT, it is not its primary objective. Instead, this inclusion highlights FIT’s broader utility as a tool for model control and adaptability, demonstrating that it performs on par with dedicated bias mitigation techniques while also offering the unique advantage of test-time steerability.
Metrics. We define the focus accuracy for a focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ as the proportion of samples where the model’s prediction aligns with the focus label, $y_{\text{focus}}$, as specified in Equation 2. Specifically, for each sample $(x,y)\in\mathcal{D}$, the model produces a prediction $\hat{y}\sim p_{\theta}(y\mid I,I_{\text{focus}},x)$ based on a fixed focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. The focus label, $y_{\text{focus}}=y_{\text{focus}}(x,y,I_{\text{focus}})$, corresponds to the target output given the focus instruction for the input $x$ with ground truth label $y$. Focus accuracy for focus instruction $I_{\text{focus}}$, denoted $\mathcal{A}_{\text{focus}}(I_{\text{focus}})$, is computed as the fraction of correct predictions with respect to the focus label:
@@
We report focus accuracy for each model on all dataset splits, using the prompt types and focus instructions detailed in Section A.4. Generations are evaluated through simple pattern matching due to the use of constrained beam decoding. Further details are provided in Section A.3.
Models and training settings. We evaluate FIT using three popular LLMs that span a range of model sizes: Llama-3.1-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.3  (Jiang et al., 2023), and Vicuna-13B-v1.5  (Chiang et al., 2023). The models are fine-tuned using parameter-efficient SFT with LoRA (Hu et al., 2021), leveraging Hugging Face’s SFTTrainer (Wolf et al., 2020) with default hyperparameters. Early stopping is applied based on validation loss, as defined in Equation 3. For generation, we use constrained beam decoding (Anderson et al., 2017) and use fully verbalised (natural language) labels during both training and testing, except for the multi-choice BBQ dataset. For further training details, refer to Section A.1.
Baselines. We compare against in the main section of the paper: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline, $\text{SFT}(y_{\text{focus}})$, follows the same setup as the FIT method (trained on sampled inputs and focus labels), but without the inclusion of focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions in FIT. The few-shot baseline involves using 5 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. In Section A.6, we include two additional baselines: zero-shot and vanilla SFT for a more complete comparison with FIT. Further details of baselines and their results in comparison to FIT can be found in Section A.5.
@@ 4.1 Validation of FIT on the SS dataset
Figure 3: SS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the SS dataset. Here, $C$ refers to the core feature (sentiment) and $S$ the spurious feature (presence of one of the keywords of “Bayesian” and “Pineapple”).
@@
Results. Figure 3 shows the focus accuracy results of the three LLMs on the SS dataset after SFT and after FIT. We see that across all focus instructions and all models, FIT shows significant improvement over the baselines, achieving very high focus accuracy.
Key takeaways. High focus accuracy on SS indicates that FIT training successfully allows the model to alter its response by considering the feature on which it is instructed to focus or not focus. This shows that the model’s behaviour can be effectively steered using FIT.
@@
Moreover, for SMNLI, we hold out specific genres at test time to evaluate our model’s ability to generalise under distribution shift when feature values change. We do this by sub-sampling a held-out portion of the MNLI dataset. During training, we use three selected genres (government, fiction, and travel) to evaluate our models. At test time, we add an additional three held-out genres (facetoface, nineeleven, and verbatim). We again ensure that $\rho_{\text{spurious}}$ is constant within each dataset split for all feature values, and use the same set of corresponding $\rho_{\text{spurious}}$ as within the SMNLI test sets described above. Further details of the SMNLI dataset can be found in Section A.11.
Figure 4: SMNLI focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the SMNLI dataset. Here, $C$ refers to the core feature (logical relationship between premise and hypothesis) and $S$ the spurious feature (genre of the underlying text)
Results. Figure 4 depicts the focus accuracy results of the three models on the SMNLI test splits. We observe that even for the more complex feature of genre, FIT achieves very high focus accuracy, significantly improving over the baselines. This demonstrates that FIT effectively trains the model to handle more complex features, allowing it to dynamically focus on or disregard these features when making predictions.
Figure 5 shows the focus accuracy of models on the distribution-shifted test sets across different values of $\rho_{\text{spurious}}$. When focusing on the core feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the second row of Figure 5). Note that, while we observe low focus accuracy when focusing on spurious features, this is expected, as the spurious labels associated with these new genres were not encountered during training. Thus when focusing on these features the model does not know what label to predict. This result highlights that the focus-tuned models have indeed learned spurious associations during training and correctly reproduces them when instructed to focus on these spurious features, even for new spurious feature values. When instructed to focus on the core feature (or even just to ignore the spurious feature), the model still shows strong generalisation in the presence of distribution shift.
Key takeaways. FIT achieves high focus accuracy on more complex features and maintains strong performance under distribution shift in terms of feature values. This demonstrates FIT’s ability to generalise to new contexts and reliably handle changing feature values, which is crucial for ensuring consistent and robust model performance in dynamic settings.
Figure 5: SMNLI focus accuracies ($\uparrow$) under distribution shift . Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on SMNLI, evaluated on distribution-shifted test sets with different feature values (genres) from training.
@@ 4.3 FIT steers behaviour in the presence of social bias data and generalises to unseen features
Figure 6: BBQ focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of models after SFT and FIT on the BBQ dataset. We include focus accuracy evaluated on bias features seen at training time (purple) and on held-out bias features seen only at test time (green).
@@ 5 Ablation
Generalisation to different test-time prompt formats. As observed in the IT literature, instruction-tuned models sometimes memorise instruction formats and struggle to follow paraphrased instructions at test time (Ghosh et al., 2024). In Section A.7 (Figure 10), we compare the performance of models on the SMNLI dataset when using the same focus instructions at training and test time versus using paraphrased instructions at test time. We generate 10 different test-time focus instructions of each instruction type defined in Equation 1 by paraphrasing the existing focus instruction using ChatGPT (OpenAI, 2022). The results show minimal variation in focus accuracy across different dataset splits and focus features, even when testing on paraphrased prompts, indicating that FIT indeed teaches models a general capacity to focus on or ignore features regardless of the specific way that focus instructions are phrased.
6 Conclusions
@@
We recommend future work explore FIT’s effectiveness in a broader range of tasks, including open-ended generation tasks like summarization or translation, which were not covered here due to the additional costs and challenges associated with curating high-quality datasets for open-ended NLG tasks. Another direction is investigating FIT’s ability to generalise across task categories.
Ethical Considerations
@@
Appendix A Appendix
A.1 FT Training and Optimisation Settings
FT Optimisation. Algorithm 1 gives precise details on how we implement FIT in practice when performing SFT of a model on a given training set. In particular, it shows how we approach optimising the FIT training objective given in Equation 3.
@@ A.1 FT Training and Optimisation Settings
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, The feature set contains $\mathcal{F}$, instruction $I$, model parameters $\theta$, batch size $B$, number of epochs $E$, step size $\eta$, and mapping function $y_{\text{focus}}=y_{\text{focus}}(x,y,I_{\text{focus}})$.
@@
6:           Sample focus instruction $I_{\text{focus}}^{b}\sim p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})$
7:           Compute $y_{\text{focus}}^{b}=y_{\text{focus}}(x^{b},y^{b},I_{\text{focus}}^{b})$
8:        end for
9:        Compute average loss given through empirical estimator of the loss defined in Equation 3 over the batch:
$L(\theta)=\frac{1}{B}\sum_{b=1}^{B}-\log p_{\theta}(y_{\text{focus}}^{b}|I,I_{\text{focus}}^{b},x^{b})$
10:        Update model parameters $\theta$ using optimiser:
$\theta\leftarrow\theta-\eta\nabla_{\theta}L(\theta)$
11:     end for
@@
13:  Output: Optimised model parameters $\theta$
@@
We implement early stopping on a held-out validation set based on the cross-entropy loss over focus labels $y_{\text{focus}}$ corresponding to randomly sampled focus instructions - this matches the context in which the models will be evaluated. We obtain this set by splitting our training set in an 80/20% for training and validation. We use a patience of $4$ validation evaluation steps, which occur after a fixed number of steps.
@@
Choice of $\rho_{\text{spurious}}$ during training. Ensuring the independence $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training is critical for inducing steerability through feature specification in the FIT framework.
@@
  • Independence $Y\perp\!\!\!\perp S$: prevents the model from leveraging the spurious feature $S$ to predict the ground-truth label $Y$. With no predictive signal flowing from $S$ to $Y$, the model must rely exclusively on the core feature $C$ for accurate label predictions. This design choice safeguards the model from overfitting to spurious correlations, thereby maintaining robust performance under distribution shifts. Moreover, removing any inherent relationship between $S$ and $Y$ ensures that for focus instruction intending for the model to utilise the core feature $C$ only during inference, the model cannot exploit a potential shortcut using $S$; it must utilise the core feature alone for prediction in this scenario enabling prediction only through the specified feature indicated through the focus instruction passed to the model.
@@
  • Independence $Y_{S}\perp\!\!\!\perp C$: serves a complementary role of preventing the model from exploiting the core feature $C$ when predicting spurious labels $Y_{S}$. By ensuring $C$ carries no information about $Y_{S}$, the model cannot use the true task feature $C$ as a shortcut for spurious-label predictions; it must again learn to only use the specified feature within the passed focus instructions alone for making predictions.
When either independence condition is violated (e.g., $Y\mathrel{\ooalign{$\perp\!\!\!\perp$\cr\hss$/$\hss}}S$ or $Y_{S}\perp\!\!\!\perp C$), the model can exploit correlations that undermine the FIT objective. For instance, if $Y\mathrel{\ooalign{$\perp\!\!\!\perp$\cr\hss$/$\hss}}S$, instructing the model to $\text{ignore}(S)$ would not necessarily prevent it from using $S$ to predict $Y$. This failure to follow instructions defeats the purpose of feature-based steerability. Thus, both $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ are necessary for building a training environment in which the model learns to precisely follow focus instructions and adjust its behaviour in accordance with specified features.
@@
To achieve this independence in our synthetic SS and SMNLI datasets, we set $\rho_{\text{spurious}}=1/N$, where $N$ is the number of class labels. Additionally, we enforce a balanced label distribution in the training set to eliminate any indirect biases that could correlate $S$ with $Y$. As shown in Section A.9 and Section A.11, these conditions are sufficient to guarantee $Y\perp\!\!\!\perp S$ in the training data, enabling the model to effectively learn steerable behaviour from focus instructions.
A.2 Limitations and Future Work
While Focus Instruction Tuning (FIT) relies on some prior knowledge of spurious features during training, this requirement aligns with existing practices in industry and research (Deldjoo, 2023; Microsoft, 2020), as well as current regulatory and ethical expectations (Guldimann et al., 2024; Zeng et al., 2024). Moreover, pre-trained models can generate focus labels automatically when explicit spurious feature labels are unavailable. For example, our method is compatible with existing and future approaches for automatically identifying and labeling spurious features (Wang & Culotta, 2020; Zhou et al., 2023; Wang et al., 2022; Zheng et al., 2024). FIT can act as a complementary mechanism, providing steerability based on identified features, whether they are manually labeled or automatically detected.
Our experiments primarily focus on classification and multiple-choice QA datasets due to the cost and challenges associated with curating high-quality datasets for open-ended NLG tasks. However, this reflects a pragmatic prioritisation of introducing a novel methodology over exhaustive data collection, rather than a limitation of FIT itself. Extending FIT to open-ended tasks, such as summarization or translation, remains an exciting direction for future research, as does exploring its ability to generalise across diverse task categories using setups similar to FLAN (Longpre et al., 2023).
Additionally, our evaluation on the HANS dataset revealed challenges when addressing overlapping or less-distinctive features. While FIT demonstrated strong performance in generalising and steering models based on identified features, overlapping heuristics can introduce ambiguity, highlighting the need for further refinements in handling such cases. Despite these limitations, FIT represents a promising foundation for enabling more robust, fair, and controllable LLMs across a range of tasks.
A.3 Evaluation Metrics

### v3

@@ 4 Experiments
In this section we empirically validate the effectiveness of FIT across a range of popular LLMs of varying sizes and on different NLP datasets, including classification and multi-choice question-answering (MCQA) tasks.
Before reporting the main results, we introduce the evaluation metric (focus accuracy) that we report, baselines, models, and training settings used throughout the experiments. In Section 4.1, we first verify that FIT performs well on the simpler SS dataset, a synthetic sentiment analysis dataset derived from SST-5 (Socher et al., 2013b). We then demonstrate in Section 4.2 that FIT generalises to more complex features and handles feature-value distribution shifts on the SMNLI dataset, a sub-sampled version of the MNLI dataset (Williams et al., 2018). Finally, in Section 4.3, we show that FIT has practical, real-world impact by effectively mitigating bias in the BBQ dataset (Parrish et al., 2022), where we further illustrate FIT’s ability to generalise to new features seen for the first time when performing inference.
Although the primary focus of FIT is on adaptively steering LLMs at inference time, which is what we focus on in this paper, we include an additional debiasing experiment for comparison on the BBQ dataset in Appendix H for completeness. While bias mitigation is a valuable and natural application of FIT, it is not its primary objective. Instead, this inclusion highlights FIT’s broader utility as a tool for model control and adaptability, demonstrating that it performs on par with dedicated bias mitigation techniques while also offering the unique advantage of test-time steerability.
Metrics. We define the focus accuracy for a focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ as the proportion of samples where the model’s prediction aligns with the focus label, $y_{\text{focus}}$, as specified in Equation 3. Specifically, for each sample $(x,y)\in\mathcal{D}$, the model produces a prediction $\hat{y}\sim p_{\theta}(\cdot\mid I,I_{\text{focus}},x)$ based on a fixed focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. The focus label, $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},s,y)$, corresponds to the target output given the focus instruction for the input $x$ with ground truth label $y$, a spurious feature value $s$ present in $x$. Focus accuracy for focus instruction $I_{\text{focus}}$, denoted $\mathcal{A}_{\text{focus}}(I_{\text{focus}})$, is computed as the fraction of correct predictions with respect to the focus label:
@@
We report focus accuracy for each model on all dataset splits, using the prompt types and focus instructions detailed in Appendix E. Generations are evaluated through simple pattern matching due to the use of constrained beam decoding (Anderson et al., 2017). Further details are provided in Appendix D.
Models and training settings. We evaluate FIT using three popular LLMs that span a range of model sizes: Llama-3.1-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.3  (Jiang et al., 2023), and Vicuna-13B-v1.5  (Chiang et al., 2023). The models are fine-tuned using parameter-efficient SFT with LoRA (Hu et al., 2021), leveraging Hugging Face’s SFTTrainer (Wolf et al., 2020) with default hyperparameters. Early stopping is applied based on validation loss, as defined in Equation 8. For generation, we use constrained beam decoding (Anderson et al., 2017) and use fully verbalised (natural language) labels during both training and testing, except for the multi-choice BBQ dataset. For further training details, refer to Appendix C.
Figure 3: SS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the SS dataset.
Baselines. We compare against the following baselines in the main section of the paper: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline, $\text{SFT}(y_{\text{focus}})$, follows the same setup as the FIT method (trained on sampled inputs and focus labels), but without the inclusion of focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions in FIT. The few-shot baseline involves using 5 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. In Appendix F, we include two additional baselines: zero-shot and vanilla SFT for a more complete comparison with FIT. Further details of baselines and their results in comparison to FIT can be found in Appendix F.
@@ 4.1 Validation of FIT on the SS dataset
Figure 4: SMNLI focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT models on the (a) SMNLI standard test sets $\mathcal{D}$, and (b) SMNLI test sets under feature value shift $\mathcal{D}^{s}$.
@@
Results. Figure 3 (a) shows the focus accuracy results of the three LLMs on the SS dataset using few-shot prompting, after $\text{SFT}(y_{\text{focus}})$ and after FIT. We see that across all focus instructions and all models, FIT shows significant improvement over the baselines, achieving very high focus accuracy across all focus instruction types and across all test sets with varying predictivity levels.
Key takeaways. High focus accuracy on SS indicates that FIT successfully steers model responses based on the feature on which it is instructed to focus or to not focus on.
@@
Moreover, for SMNLI, we hold out specific genres at test time to evaluate our model’s ability to generalise under distribution shift when feature values change. We do this by sub-sampling a held-out portion of the MNLI dataset. During training, we use three selected genres (government, fiction, and travel) to evaluate our models. At test time, we additionally add three held-out genres (facetoface, nineeleven, and verbatim). We again ensure that $\rho_{\text{spurious}}$ is constant within each dataset split for all feature values, and use the same set of corresponding $\rho_{\text{spurious}}$ as within the SMNLI test sets described above. Further details of the SMNLI dataset can be found in Appendix K.
Results. Figure 4 (a) depicts the focus accuracy results of the three models on the SMNLI test splits. We observe that for the more complex feature of genre, FIT achieves very high focus accuracy, significantly improving over the baselines. This demonstrates that FIT effectively trains the model to handle more complex features, allowing it to dynamically focus on or disregard these features when making predictions.
Figure 4 (b) shows the focus accuracy of models on the feature-shifted test sets. When focusing on the core feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the third row of Figure 4) (b). Note that, while we observe low focus accuracy when focusing on spurious features, this is expected, as the spurious labels associated with these new genres were not encountered during training. Thus when focusing on these features the model does not know what label to predict. This result highlights that the focus-tuned models have indeed learned spurious associations during training and correctly reproduces them when instructed to focus on these spurious features, even for new spurious feature values. When instructed to focus on the core feature or when instructed to ignore the spurious feature, the model still shows strong generalisation in the presence of distribution shift.
Key takeaways. FIT achieves high focus accuracy on more complex features and maintains strong performance under distribution shift in terms of feature values. This demonstrates FIT’s ability to generalise to new contexts and reliably handle changing feature values.
@@ 4.3 FIT steers behaviour in the presence of social bias data and generalises to unseen features
Figure 5: BBQ focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the BBQ dataset.
@@ 5 Ablation
Generalisation to different test-time prompt formats. As observed in the IT literature, instruction-tuned models sometimes memorise instruction formats and struggle to follow paraphrased instructions at test time (Ghosh et al., 2024). In Appendix G (Figure 9), we compare the performance of models on the SMNLI dataset when using the same focus instructions at training and test time versus using paraphrased instructions at test time. We generate 10 different test-time focus instructions of each instruction type defined in Equation 2 by paraphrasing the existing focus instruction using ChatGPT (OpenAI, 2022). The results show minimal variation in focus accuracy across different dataset splits and focus features, even when testing on paraphrased prompts, indicating that FIT indeed teaches models a general capacity to focus on or ignore features regardless of the specific way that focus instructions are phrased.
Instruction Following After FIT. Prior studies suggest that SFT can impair the instruction-following abilities of LLMs (Fu et al., 2024; Dou et al., 2024). We evaluate whether FIT impacts instruction adherence by comparing pre-trained and FIT models trained on the SMNLI dataset (see Section 4.2). Using 500 samples from the Alpaca-GPT instruction-tuning dataset (Peng et al., 2023), responses were rated by GPT-4o (Achiam et al., 2023) on a 1–5 scale, where 5 denotes perfect alignment and 1 indicates a completely incorrect response.
LLaMA Mistral Vicuna
Pre-Trained Avg. Rating 3.51 3.65 3.46
FIT Avg. Rating 3.45 3.65 3.50
$p$-value $\bm{0.57}_{>0.05}$ $\bm{0.81}_{>0.05}$ $\bm{0.41}_{>0.05}$
Table 1: Instruction-Following Results. For each model (columns), we report the pre-trained and FIT average GPT-40 ratings, and the two-sided Wilcoxon Signed-Rank $p$-value testing the difference between the distributions of ratings.
To test for significant differences in instruction-following performance, we conducted a two-sided Wilcoxon Signed-Rank Test (Wilcoxon, 1992) on paired ratings. The null hypothesis assumes no difference in the median ratings between pre-trained and FIT models. The results, summarized in Table 1, show no statistically significant differences ($p>0.05$) in instruction-following performance across all models. These findings demonstrate that FIT maintains general instruction-following capabilities whilst additionally enhancing test-time steerability.
6 Conclusion
@@
Appendix A Impact Statement
@@
Appendix B Limitations and Future Work
Requirement for annotated spurious features. While FIT relies on prior identification of spurious features and their focus labels, this requirement does not limit its practical applicability. Instead, it reflects standard industry and research practices for constructing transparent and reliable models. Below, we clarify how FIT remains adaptive and versatile even when feature annotation is partial or evolving:
  • Alignment with Established Practices: FIT’s reliance on pre-identified spurious features aligns with widely adopted industry and research norms (OpenAI, 2024; Microsoft, 2020). Identifying potential spurious features and confounders in datasets is a foundational step in achieving robust machine learning systems. This process ensures that both training and validation phases are informed by an understanding of data correlations, minimising the risk of deploying models with unknown biases.
  • Regulatory and Ethical Expectations: Regulatory frameworks and ethical guidelines increasingly require the explicit identification and mitigation of problematic features (of the European Parliament, 2016) Corresponding initiatives aim to define and enforce measurable categories of “violating behaviour” in AI models. By providing a mechanism to steer model behaviour based on these identified features, FIT effectively complements efforts to promote fair and transparent predictions (Guldimann et al., 2024; Zeng et al., 2024).
  • Post-Deployment Mitigation: Despite careful pre-deployment analysis, spurious features or correlations may only become apparent once a model is in active use. FIT accommodates this by allowing developers to incorporate newly identified spurious features via updated focus instructions, enabling rapid iterative refinement without retraining from scratch. This adaptability ensures continuous improvement, even in highly dynamic environments.
  • FIT’s Versatility Without Exhaustive Pre-Identification: Crucially, FIT does not require an exhaustive list of spurious features to be effective. For instance, a user can provide focus instructions such as “focus on casual” without enumerating every possible irrelevant attribute in the dataset. This flexibility expands FIT’s applicability to scenarios where feature annotation is incomplete or ongoing.
  • Compatibility with Automated Spurious Feature Identification: FIT also works seamlessly with automated methods for detecting spurious features (Wang & Culotta, 2020; Wang et al., 2022; Zhou et al., 2023; Zheng et al., 2024). Whether spurious features are labelled manually or derived from algorithmic detection, they can be harnessed by FIT’s focus instructions at inference time. This compatibility enables a comprehensive approach to managing known issues and responding to newly uncovered features as they arise.
In summary, annotating spurious features beforehand is not a strict limitation. FIT can be flexibly applied, allowing model behaviour to evolve in tandem with new feature discoveries or changing requirements, making it a broadly applicable technique for steering model outputs based on both prior knowledge and ongoing insights.
Scope of Experiments and Extensions to Open-Ended Tasks. Our experiments primarily focus on classification and multiple-choice QA datasets due to the cost and challenges associated with curating high-quality datasets for open-ended NLG tasks. However, this reflects a pragmatic prioritisation of introducing a novel methodology over exhaustive data collection, rather than a limitation of FIT itself. Extending FIT to open-ended tasks, such as summarisation or translation, remains an exciting direction for future research, as does exploring its ability to generalise across diverse task categories using setups similar to FLAN (Longpre et al., 2023).
Overlapping Features and Ambiguities. Additionally, our evaluation on the HANS dataset Appendix L revealed challenges when addressing overlapping or less-distinctive features. While FIT demonstrated strong performance in generalising and steering models based on identified features, overlapping heuristics can introduce ambiguity, highlighting the need for further refinements in handling such cases. Despite these limitations, FIT represents a promising foundation for enabling more robust, fair, and controllable LLMs across a range of tasks.
Appendix C FT Training and Optimisation Settings
FT Optimisation. Algorithm 1 gives precise details on how we implement FIT in practice when performing ERM of a model on a given training set. In particular, it shows how we approach optimising the FIT training objective given in Equation 8.
@@ Appendix C FT Training and Optimisation Settings
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, The feature set contains $\mathcal{F}$, instruction $I$, model parameters $\theta$, batch size $B$, number of epochs $E$, step size $\eta$, and mapping function $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},y,s)$.
@@
6:           Identify spurious feature value $s^{b}$ in $x^{b}$.
7:           Sample focus instruction $I_{\text{focus}}^{b}\sim p_{\mathcal{I}_{\text{focus}}}$,
8:           Compute $y_{\text{focus}}^{b}=y_{\text{focus}}(I_{\text{focus}}^{b},s^{b},y^{b})$
9:        end for
10:        Compute average loss given through empirical estimator of the loss defined in Equation 8 over the batch:
$\ell(\theta)=\frac{1}{B}\sum_{b=1}^{B}-\log p_{\theta}(y_{\text{focus}}^{b}|I,I_{\text{focus}}^{b},x^{b})$
11:        Update model parameters $\theta$ using optimiser:
$\theta\leftarrow\theta-\eta\nabla_{\theta}\ell(\theta)$
@@
13:  end for
14:  Output: Optimised model parameters $\theta$
@@
We implement early stopping on a held-out validation set based on the cross-entropy loss over focus labels $y_{\text{focus}}$ corresponding to randomly sampled focus instructions - this matches the context in which the models will be evaluated. We obtain this set by splitting our training set in a 90/10% ratio for training and validation splits respectively. We use a patience of $4$ validation evaluation steps, which occur after a fixed number of steps.
@@
Choice of $\rho_{\text{spurious}}$ during training. In our synthetic experiments, we set up a controlled environment by imposing two independence conditions: $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$. These ensure that (i) the ground-truth label cannot be predicted using the spurious feature $S$, and (ii) the spurious label cannot be predicted using the core feature $C$. By removing direct correlations between these features and labels, the model is leans to focus on the specified feature, without being influenced by the other feature, avoiding any potential shortcuts that could be exploited if these conditions did not hold.
@@
  • Independence $Y\perp\!\!\!\perp S$: This condition prevents the model from leveraging spurious feature $S$ to predict ground-truth label $Y$. With no predictive signal from $S$ to $Y$, the model must rely exclusively on the core feature $C$ for accurate label predictions. This design choice safeguards the model from overfitting to spurious correlations, thereby maintaining robust performance under distribution shifts. Moreover, removing any inherent relationship between $S$ and $Y$ ensures that for focus instruction intending for the model to utilise the core feature $C$ only during inference, the model cannot exploit a potential shortcut using $S$; it must utilise the core feature alone for prediction in this scenario enabling prediction only through the specified feature indicated through the focus instruction passed to the model.
@@
  • Independence $Y_{S}\perp\!\!\!\perp C$: This condition serves a complementary role in preventing the model from exploiting the core feature $C$ when predicting spurious labels $Y_{S}$. By ensuring $C$ carries no information about $Y_{S}$, the model cannot use the true task feature $C$ as a shortcut for spurious-label predictions; it must again learn to only use the specified feature within the passed focus instructions alone for making predictions.
While these conditions represent an ideal setting, they are not strictly necessary for FIT to work in practice. Indeed, real-world data rarely satisfies such perfect independence, and we illustrate the robustness of the method in more realistic scenarios through our BBQ experiments in Section 4.3 where correlations between $Y$ and $S$ or between $C$ and $Y_{S}$ may exist as no subsampling or dataset manipulations have been made. By examining both the controlled environment and more naturalistic datasets, we demonstrate that our approach can handle scenarios with varying degrees of spurious correlations.
@@
To achieve this independence in our synthetic SS and SMNLI datasets, we set $\rho_{\text{spurious}}=1/N$, where $N$ is the number of class labels. Additionally, we enforce a balanced label distribution in the training set to eliminate any indirect biases that could correlate $S$ with $Y$. As shown in Appendix I and Appendix K, these conditions are sufficient to guarantee $Y\perp\!\!\!\perp S$ in the training data, enabling the model to effectively learn steerable behaviour from focus instructions.
Appendix D Evaluation Metrics